OpenStudy (anonymous):

6.03] What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 4 −2 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (foreverandalways13):

do you know elimination?

OpenStudy (anonymous):

no

OpenStudy (foreverandalways13):

ok then i'll try to explain it to u...

OpenStudy (foreverandalways13):

2(3x+2y=8) 3(2x+3y=2) multiply that out then tell me what u get...

OpenStudy (foreverandalways13):

hi uri...

OpenStudy (foreverandalways13):

@Kylo744

OpenStudy (anonymous):

had to fined something to write with

OpenStudy (foreverandalways13):

oh ok...lol

OpenStudy (foreverandalways13):

yup! thanks uri:)

OpenStudy (uri):

waittt

OpenStudy (foreverandalways13):

k..

OpenStudy (anonymous):

6x+4y=16 6x+9y=9

OpenStudy (foreverandalways13):

ok now the x cancel...then subtract the 4 and 9 and the 16 and 9...then divide your answer and u get y...

OpenStudy (foreverandalways13):

then u just plug y in for one of the ORIGINAL equations and solve for x... does that help?

OpenStudy (anonymous):

x can be any number as there are infinitely many solutions to this system

OpenStudy (foreverandalways13):

yup! GOOD JOB!!

OpenStudy (anonymous):

thank u for your help

OpenStudy (foreverandalways13):

np....let me know if u need help with something else..

OpenStudy (anonymous):

6.02] Carlos solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x + 10 + 4x = 10 5x + 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He did not distribute correctly He combined like terms incorrectly, it should have been 4x instead of 5x He made no mistake

OpenStudy (foreverandalways13):

give me like 5 seconds:) I actually need to write this one out:D

OpenStudy (foreverandalways13):

he didn't make a mistake...trick question:)

OpenStudy (foreverandalways13):

waiiittt!

OpenStudy (foreverandalways13):

I don't think he distributed properly...he used substitution so im not sure but I think that's right:)

OpenStudy (anonymous):

He did not distribute correctly

OpenStudy (foreverandalways13):

ok..cool! I was right:)

OpenStudy (foreverandalways13):

im leah btw!

OpenStudy (foreverandalways13):

just message me if anything else...lol

OpenStudy (anonymous):

Help me with one more

OpenStudy (foreverandalways13):

ok

OpenStudy (anonymous):

6.03] If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? x − y = 6 4x − 3y = 12 Add the two equations together Multiply the first equation by −4 Multiply the first equation by 3 Multiply the first equation by −3

OpenStudy (foreverandalways13):

whats 4+-3?

OpenStudy (anonymous):

1

OpenStudy (foreverandalways13):

thanks for fanning me btw:) oh yeah duh:)

OpenStudy (foreverandalways13):

the second one...I had to work it out first:)

OpenStudy (anonymous):

ok thank u

OpenStudy (foreverandalways13):

np! I had fun...im like into way more complicated stuff soooo this was a relaxing thing:)

OpenStudy (foreverandalways13):

welcome to OS!!!!